# TIGER: A Generating-Then-Ranking Framework for Practical Python Type Inference

### v1

@@ V-B4 Comparison with Generation-based Approaches
Among the three generation-based baseline approaches, CodeT5-zs exhibits poor accuracy due to its application in a zero-shot setting, while the fine-tuned CodeT5-ft and TypeGen achieve the second-best or the co-best results for specific type categories and metrics. Specifically, CodeT5-ft demonstrates effectiveness for elementary types (Column “Ele”) and generic types (Column “Gen”), but is relatively ineffective for user-defined types (Column “Usr (Unseen)”); whereas TypeGen achieves good effectiveness for user-defined types but struggles with handling generic types effectively.
CodeT5-ft’s lower efficacy in handling user-defined types stems from its inability to grasp project-specific contexts, while the limited effectiveness of TypeGen for generic and unseen user-defined types can be attributed to its reliance on prompt engineering techniques to leverage the capability of LLMs, which often poses challenges in designing an optimal prompt template [18]. In comparison, our approach, TIGER, consistently achieves substantial inference accuracies for all type categories, especially (unseen) user-defined types (Column “Usr (Unseen)”), with enhancements of 16.5% (25.6%) and 11.2% (20.1%) compared to CodeT5-ft and TypeGen, respectively, in Top-5 Exact Match. These improvements are attributed to the comprehensive visible candidate pool and robust inference capacity integrated within the two-stage inference framework. Considering the prevalence of encountering (unseen) user-defined types in real-world projects, our approach proves to be more practical and applicable.
@@ VII Conclusion and Future Work
In the future, we plan to extend the applicability of TIGER to other dynamic programming languages such as JavaScript. Additionally, we aim to adapt TIGER to address other practical type inference challenges, including those related to partial Python functions.
VIII Data Availability

### v3

@@ V-B4 Comparison with Generation-based Approaches
Among the three generation-based baseline approaches, CodeT5-zs exhibits poor accuracy due to its application in a zero-shot setting, while the fine-tuned CodeT5-ft and TypeGen achieve the second-best or the co-best results for specific type categories and metrics. Specifically, CodeT5-ft demonstrates effectiveness for elementary types (Column “Ele”) and parameterized types (Column “Par”), but is relatively ineffective for user-defined types (Column “Usr (Unseen)”); whereas TypeGen achieves good effectiveness for user-defined types but struggles with handling parameterized types effectively.
CodeT5-ft’s lower efficacy in handling user-defined types stems from its inability to grasp project-specific contexts, while the limited effectiveness of TypeGen for parameterized and unseen user-defined types can be attributed to its reliance on prompt engineering techniques to leverage the capability of LLMs, which often poses challenges in designing an optimal prompt template [18]. In comparison, our approach, TIGER, consistently achieves substantial inference accuracies for all type categories, especially (unseen) user-defined types (Column “Usr (Unseen)”), with enhancements of 16.5% (25.6%) and 11.2% (20.1%) compared to CodeT5-ft and TypeGen, respectively, in Top-5 Exact Match. These improvements are attributed to the comprehensive visible candidate pool and robust inference capacity integrated within the two-stage inference framework. Considering the prevalence of encountering (unseen) user-defined types in real-world projects, our approach proves to be more practical and applicable.
@@ V-E Ablation Study (RQ4)
V-F Case Study
We have identified certain suboptimal cases, as illustrated in Figure 6, concerning user-defined types.
(a) Example 1: Predictions for Missing Type “promise
(b) Example 2: Predictions for Missing Type “Data
Figure 6: Two Suboptimal Cases
In Figure 6(a), TIGER’s generation model produces complex parameterized types for a type-missed function, with their corresponding generative likelihoods (i.e., $lik$). However, the expected type “promise” does not appear in the candidate list. Conversely, TIGER’s similarity model ranks the ground-truth type as the top-1 candidate with a high similarity score (i.e., $sim$). When the two models are combined, TIGER eventually returns “promise” as the 2nd result, just after “Thenable”. Similarly, in Figure 6(b), the expected “Data” type is absent from the candidates generated by the generation model but is ranked 1st by the similarity model. However, in the final results, it only appears in 5th place.
These suboptimal outcomes can be attributed to the fact that some user-defined types, which are not generated as candidates by the generation model, tend to have lower likelihoods compared to model-generated candidates. As a result, even with high similarities from the similarity model, their final scores, as calculated by Eq. 6, remain lower. While we have attempted to adjust the weights of the two measures in Eq. 6, finding an optimal balance has proven challenging. In the future, we plan to explore more effective methods for combining these two measures, such as incorporating likelihood into the contrastive loss.
@@ VII Conclusion and Future Work
In the future, we plan to extend the applicability of TIGER to other dynamic programming languages such as JavaScript and explore the further integration with static analysis for type validation. Additionally, we aim to adapt TIGER to address other practical type inference challenges, including those related to partial Python functions.
Data Availability
@@
Acknowledgement
This research / project is supported by the National Research Foundation, Singapore, and the Cyber Security Agency under its National Cybersecurity R&D Programme (NCRP25-P04-TAICeN). Any opinions, findings and conclusions or recommendations expressed in this material are those of the author(s) and do not reflect the views of National Research Foundation, Singapore and Cyber Security Agency of Singapore.